# Three-Stream Temporal-Shift Attention Network Based on Self-Knowledge Distillation for Micro-Expression Recognition

### v2

@@ 5 Conclusion
In this paper, we propose a three-stream temporal-shift attention network based on self-knowledge distillation (SKD-TSTSAN) for ME recognition. First, we employ learning-based motion magnification modules to enhance muscle motion to make the ME features more discriminative. To suppress redundant information in the local-spatial stream and focus on important facial regions, we introduce a channel attention mechanism using the ECA modules. In addition, TSMs are incorporated into the dynamic-temporal stream for temporal modeling through information exchange in the temporal dimension. Furthermore, we introduce SKD into ME recognition by adding auxiliary classifiers into the network, taking the deepest part of the network as the teacher model for the shallower part and further improving the performance. Finally, extensive experiments on the CASME II, SAMM, MMEW, and CAS(ME)3 datasets show that our SKD-TSTSAN outperforms other cutting-edge MER methods.
In this paper, we classify MEs into a set of emotion categories using the onset frames, apex frames, and offset frames labeled in the dataset. However, to achieve an end-to-end ME analysis system, we still need to design an ME spotting method to identify key-frames. Therefore, in future work, we
will explore ME spotting methods in long videos. Additionally, many current research efforts treat the ME spotting and recognition tasks separately (Song et al.,, 2019). However, these two tasks are strongly related and both require extracting subtle facial muscle motion features. Therefore, we will attempt to integrate ME spotting with our designed ME recognition network to implement ME analysis in a unified network.

### v3

@@ 5 Conclusion
In this paper, we propose a three-stream temporal-shift attention network based on self-knowledge distillation (SKD-TSTSAN) for ME recognition. First, we employ learning-based motion magnification modules to enhance muscle motion to make the ME features more discriminative. To suppress redundant information in the local-spatial stream and focus on important facial regions, we introduce a channel attention mechanism using the ECA modules. In addition, TSMs are incorporated into the dynamic-temporal stream for temporal modeling through information exchange in the temporal dimension. Furthermore, we introduce SKD into ME recognition by adding auxiliary classifiers into the network, taking the deepest part of the network as the teacher model for the shallower part and further improving the performance. Finally, extensive experiments on the CASME II, SAMM, SMIC, MMEW, and CAS(ME)3 datasets show that our SKD-TSTSAN outperforms other cutting-edge MER methods.
In this paper, we classify MEs into a set of emotion categories using the onset frames, apex frames, and offset frames labeled in the dataset. However, to achieve an end-to-end ME analysis system, we still need to design an ME spotting method to identify key-frames. Therefore, in future work, we will explore ME spotting methods in long videos. Additionally, many current research efforts treat the ME spotting and recognition tasks separately (Song et al.,, 2019). However, these two tasks are strongly related and both require extracting subtle facial muscle motion features. Therefore, we will attempt to integrate ME spotting with our designed ME recognition network to implement ME analysis in a unified network.